# ContextFlow: Context-Aware Flow Matching for Trajectory Inference from Spatial Omics Data

### v1

@@ 4.1 Axolotl Brain Regeneration
(a) Brain Regeneration
(b) Organogenesis
Figure 2: KL-Divergence between predicted and ground-truth cell type distributions.
@@ 4.2 Mouse Embryo Organogenesis
Figure 2 reports the KL-Divergence between normalized histograms of predicted and ground-truth cell types from ContextFlow and MOTFM. In both cases, CTF exhibits lower divergence on average across time points, indicating that the trajectories generated by our model better preserve the biological composition of cell types over time. The cell type progression is further visualized in Figure 9 (Appendix H.6). We show that ContextFlow predicts temporal cell type trajectories that evolve smoothly and consistently across consecutive developmental stages. Early progenitor populations, such as neural crest and mesenchyme, progressively diminish as development advances, while terminal fates, including muscle, cartilage primordium, and liver, emerge at later stages. Major lineages such as brain, heart, and connective tissue remain continuous throughout, demonstrating that ContextFlow captures biologically coherent and temporally consistent developmental dynamics.
@@ 4.3 Liver Regeneration
Finally, we evaluate ContextFlow on a Visium spatial transcriptomics dataset profiling the temporal dynamics of mouse liver regeneration following acetaminophen-induced injury (Ben-Moshe et al., 2021), collected across three distinct regeneration stages. Unlike the earlier datasets resolved at single-cell resolution, Visium data is captured at the level of $55$ micron diameter spots, capturing the joint expression of multiple cells. Since direct cell-type information is unavailable, we restrict evaluation to the 2-Wasserstein distance. Moreover, since evaluation is performed on the middle of the three time points, IVP and next-step predictions coincide. Table 4 presents the results. Consistent with the previous findings, CTF-H achieved the lowest reconstruction error, indicating that incorporating contextual information improves trajectory estimation even in aggregated spot-level measurements.
@@ 5 Conclusion
We introduced ContextFlow, a contextually aware flow matching framework that leverages spatial priors and biologically motivated constraints to learn more plausible trajectories from snapshot spatial transcriptomic data, addressing a central challenge of existing methods. The entropic variant of ContextFlow is theoretically grounded, which always yields OT couplings constrained by prior knowledge, promoting stability and consistency with the imposed contextual constraints. Across three diverse datasets, we showed that ContextFlow consistently improves over state-of-the-art baselines even in challenging Initial Value Problem sampling settings, underscoring the importance of our contextually informed priors. In addition, we demonstrated that our framework reduces the number of biologically implausible couplings and results in coherent and temporally consistent developmental trajectories while maintaining strong quantitative performance across Wasserstein, MMD, and Energy metrics. These results highlight the value of embedding biological context into generative flow models. Future works can adapt our methods to reconstruct tissues and learn spatial latent dynamics by formulating the flow in space (rather than time), or leverage multi-marginal OT formulations for optimizing temporal flows. Looking forward, ContextFlow offers a principled foundation for modeling perturbations and disease progression, bridging generative power with biological interpretability.
@@ A.1 Flow Matching
Normalizing flows provide a parametric framework for characterizing transformations of a random variable into desired distributions (Papamakarios et al., 2021). These transformations can be realized through either finite (Rezende & Mohamed, 2015) or continuous compositions (Chen et al., 2018). The loss functions used in such formulations typically require computing Jacobians or integrating the flows at each forward pass, making them computationally expensive. Flow matching (FM) (Lipman et al., 2023; Albergo & Vanden-Eijnden, 2022; Liu et al., 2022) addresses this limitation by reducing the training of the velocity field to a regression problem, thereby making normalizing flows substantially more scalable. To ensure valid conditional paths at intermediate time points, samples are coupled either randomly or via optimal transport (Pooladian et al., 2023; Tong et al., 2024). Owing to this scalability, FM has been rapidly adopted across scientific domains, including biology and the life sciences (Li et al., 2025). In transcriptomics, for example, Klein et al. (2024) employed an FM backbone to approximate OT maps for drug response modeling and cross-modal translation tasks. Entropic OT formulations have also been applied to infer cellular trajectories (Tong et al., 2024; Rohbeck et al., 2025), generate imaging-based cell morphology changes (Zhang et al., 2025), and simulate spatial transcriptomics data from histology images (Huang et al., 2025).
Despite these advances, existing work does not address how to meaningfully incorporate biological prior knowledge to constrain the velocity field, limiting the biological plausibility of inferred trajectories.
@@ A.2 Optimal Transport
Omics studies frequently generate uncoupled measurements across conditions, modalities, or time points, which must be integrated into a unified representation to provide a more comprehensive view of the underlying biology. Optimal transport (OT) has recently gained popularity for this task, as it provides a geometry-based approach to couple probability distributions (Bunne et al., 2024; Klein et al., 2025). In spatial transcriptomics (ST), several OT formulations have been introduced depending on context. For instance, Zeira et al. (2022) and Liu et al. (2023) proposed PASTE and PASTE2 to align ST data from adjacent tissue slices, while DeST-OT (Halmos et al., 2025) integrates spatio-temporal slices by modeling cell growth and differentiation. Rahimi et al. (2024) developed DOT, a multi-objective OT framework for mapping features across scRNA-seq and spatially resolved assays, and Ceccarelli et al. (2025) introduced TOAST, a spatially regularized OT framework for slice alignment and annotation transfer.
While these methods are primarily designed to align biological data across space, time, or modality, they do not address the problem of trajectory inference toward biologically plausible solutions, leveraging biological priors to constrain or bias the transport plan.
Appendix B Proofs of Main Theoretical Results
Proposition 1.
@@ Appendix B Proofs of Main Theoretical Results
$\Pi^{*}$ $={\arg\!\min}_{\Pi\geq 0}\;\sum_{k,l}\Pi_{kl}C_{kl}+\epsilon\sum_{k,l}\Pi_{kl}\log\!\left(\Pi_{kl}\right).$
@@
From Proposition 3 Proposition 2, we have:
@@
$\Pi^{*}$ $={\arg\!\min}_{\Pi\geq 0}\;\sum_{k,l}\Pi_{kl}C_{kl}+\epsilon\sum_{k,l}\Pi_{kl}\log\!\left(\Pi_{kl}\right).$
@@
$\Pi^{*}$ $={\arg\!\min}_{\Pi\geq 0}\;\sum_{k,l}\Pi_{kl}C_{kl}+\epsilon\sum_{k,l}\Pi_{kl}\log\!\left(\Pi_{kl}\right).$
@@
$\epsilon\log\left(\frac{\Pi^{*}_{kl}}{M_{kl}}\right)=f_{k}-C_{kl}-g_{l}$
@@
From equations 16 13, 17 14, and 18 15 above, we get:
@@
Sinkhorn Algorithm uses these updates, iteratively, and these updates are shown to converge in Franklin & Lorenz (1989). Thus, Sinkhorn Algorithm can be used for the ContextFlow’s Prior Aware Entropy Regularized (PAER) (CTF-H) formulation.
From equation (9) 13, we get:
@@
Appendix C Effects of normalization on Prior aware cost matrix
From Peyré et al. (2019), we know that optimal MOTFM coupling takes the form $\Pi^{*}_{\text{EOT}}=\text{diag}(u)\cdot K\cdot\text{diag}(v)$, where $K$ is the kernel matrix such that $K]_{ij}=\exp(\frac{-c_{ij}}{\epsilon})$, with $u,v$ satisfying marginalization constraints $u\odot Kv=a$ and $K^{T}u\odot v=b$. Sinkhorn updates are given by:
@@ Appendix C Effects of normalization on Prior aware cost matrix
$\Pi^{*}$ $={\arg\!\min}_{\mathbf{\Pi}\geq 0}\;\sum_{k,l}\Pi_{kl}C_{kl}+\epsilon\sum_{k,l}\Pi_{kl}\log\!\left(\Pi_{kl}\right).$
@@
$\mathbf{\Pi}^{*}$ $={\arg\!\min}_{\mathbf{\Pi}\geq 0}\;\sum_{k,l}\Pi_{kl}C_{kl}+\epsilon\sum_{k,l}\Pi_{kl}\log\!\left(\Pi_{kl}\right)$
@@
Corollary 1 can also be interpreted as supporting the results of Proposition 1 and our intuition that normalizing has the same effect on the kernel matrix as increasing $\epsilon$, leading to more diffused couplings or couplings with increased entropy.
Appendix D ContextFlow Algorithm
@@
Algorithm 1 ContextFlow (CTF): Flow Matching with Spatial-Context-Aware OT Couplings
@@
1:Input: gene data $\{\mathbf{X}_{t_{1}},\cdots,\mathbf{X}_{t_{m+1}}\}$, spatial data $\{\mathbf{S}_{t_{1}},\ldots,\mathbf{S}_{t_{m+1}}\}$, parameters $\lambda$, $\alpha$, $\epsilon$, $\sigma$, $\eta$
2:Output: neural velocity vector field $u_{\theta}$
3:Initialize $\theta$
4:while training do
5:  for $i=1,2,\ldots,m$ do
6:   Sample a batch $\mathcal{B}=\{(\bm{x}_{i},\bm{x}_{i+1}):(\bm{x}_{i},\bm{x}_{i+1})\sim(\mathbf{X}_{t_{i}},\mathbf{X}_{t_{i+1}})\}$
7:   Construct TPM: $\mathbf{M}_{i,i+1}(\mathcal{B})\triangleright$ $\mathbf{M}_{i,i+1}$ is defined in Equation 10
8:   if “prior-aware cost matrix” then
9:     $C_{kl}\leftarrow\alpha\cdot\|\bm{x}_{i}(k)-\bm{x}_{i+1}(l)\|_{2}^{2}+(1-\alpha)\cdot[\mathbf{M}_{i,i+1}]_{kl}$ for any pair $(k,l)$
10:     $\mathbf{K}\leftarrow\exp(-\mathbf{C}/\epsilon)$
11:   else if “prior-aware entropy regularization” then
12:     $C_{kl}\leftarrow\|\bm{x}_{i}(k)-\bm{x}_{i+1}(l)\|_{2}^{2}$ for any pair $(k,l)$
@@
13:     $\mathbf{K}\leftarrow\widehat{\mathbf{M}}_{i,i+1}\odot\exp(-\mathbf{C}/\epsilon)$ $\triangleright$ $\widehat{\mathbf{M}}_{i,i+1}$ is defined in Equation 12
14:   end if
15:   Initialize $\bm{a}\leftarrow\frac{1}{n_{i}}\mathbf{1}_{n_{i}}$, $\bm{b}\leftarrow\frac{1}{n_{i+1}}\mathbf{1}_{n_{i+1}}$, $\bm{u}\leftarrow\mathbf{1}_{n_{i}}$, $\bm{v}\leftarrow\mathbf{1}_{n_{i+1}}$
16:   while not converged do
17:     $\bm{u}\leftarrow\bm{a}\oslash(\mathbf{K}\bm{v}),\>\bm{v}\leftarrow\bm{b}\oslash(\mathbf{K}^{\top}\bm{u})$ $\triangleright$ Run Sinkhorn algorithm
18:   end while

### v2

@@ 4 Experiments
Sampling. A trained velocity field can be evaluated through the samples it generates. We consider two variants of sampling configurations: initial value problem sampling (IVP) and next-step sampling (Next Step). In particular, IVP integrates the learned gradients from the first observed batch of cells and evolves them to a later time point. IVP provides the most comprehensive evaluation of flow quality, as errors can accumulate across steps. In contrast, Next Step integrates the gradient only from the most recently observed batch of cells, thus limiting error propagation but providing a less stringent test of long-term trajectory fidelity.
4.1 Axolotl Brain Regeneration
We first evaluate ContextFlow on longitudinal Stereo-seq spatial transcriptomic data coming from a post-traumatic brain regeneration study of the Salamander (axolotl telencephalon) species (Wei et al., 2022). The dataset contains samples from five developmental stages, with replicates collected from different individual organisms at each stage. For our CTF-C method, we present the best ablated $\alpha$, with full ablation results across different $\alpha$ values in Appendix H.
For extrapolation, we evaluate generation at the last holdout time point, which represents the most challenging test of generalizability for the velocity fields $u_{\theta}$, as it lies outside the training time horizon. As shown in Table 3, CTF-H again consistently achieves the best overall performance, particularly under IVP-Sampling, where errors are most likely to accumulate. This result further reinforces the robustness and reliability of CTF-H across the entire sampling horizon. Figure 4 in Appendix G.1 further illustrates that incorporating spatial priors enables ContextFlow to produce substantially fewer biologically implausible couplings compared to its context-free counterpart.
@@ 4.2 Mouse Embryo Organogenesis
(a) Brain Regeneration
(b) Organogenesis
Figure 2: Comparisons of predicted and true cell type distributions.
Figure 2 reports the KL-Divergence between normalized histograms of predicted and ground-truth cell types from ContextFlow and MOTFM. In both cases, CTF exhibits lower divergence on average across time points, indicating that the trajectories generated by our model better preserve the biological composition of cell types over time. The cell type progression is further visualized in Figure 10 (Appendix H.6). We show that ContextFlow predicts temporal trajectories of cell types that evolve smoothly and consistently across consecutive developmental stages. Early progenitor populations, such as neural crest and mesenchyme, progressively diminish as development progresses, while terminal fates, including muscle, the cartilage primordium and the liver, emerge at later stages. Major lineages such as brain, heart, and connective tissue remain continuous throughout, confirming that ContextFlow captures biologically coherent and temporally consistent developmental dynamics.
@@ 4.3 Liver Regeneration
Finally, we evaluate ContextFlow on a Visium spatial transcriptomics dataset profiling the temporal dynamics of mouse liver regeneration following acetaminophen-induced injury (Ben-Moshe et al., 2021), collected across three distinct regeneration stages. Unlike earlier datasets resolved at single-cell resolution, Visium data is captured at the level of $55$ micron diameter spots, capturing the joint expression of multiple cells. Since direct cell-type information is unavailable, we restrict evaluation to the 2-Wasserstein distance. Moreover, since evaluation is performed on the middle of the three time points, IVP and next-step predictions coincide. Table 5 summarizes the comparisons between MOTFM and our methods. Consistent with the previous findings, CTF-H achieved the lowest reconstruction error, indicating that incorporating contextual information improves trajectory estimation even in aggregated spot-level measurements.
@@ 5 Conclusion
We introduced ContextFlow, a contextually aware flow matching framework that leverages spatial priors and biologically motivated constraints to learn more plausible trajectories from snapshot spatial transcriptomic data. The entropic variant of ContextFlow is theoretically grounded, which produces OT couplings that are constrained by prior knowledge, thereby promoting stability and consistency with the imposed contextual constraints. Across three diverse datasets, we showed that ContextFlow consistently outperforms state-of-the-art baselines, even in challenging Initial Value Problem sampling settings, underscoring the importance of our contextually informed priors. In addition, we demonstrated that our framework reduces the number of biologically implausible couplings and yields coherent, temporally consistent developmental trajectories while maintaining strong quantitative performance across Wasserstein, MMD, and Energy metrics. These results highlight the value of embedding biological context into generative flow models. Future work can adapt our methods to reconstruct tissues and learn spatial latent dynamics by formulating the flow in space (rather than time), or by leveraging multi-marginal OT formulations to optimize temporal flows. Looking forward, ContextFlow offers a principled foundation for modeling perturbations and disease progression, potentially bridging generative power with biological interpretability.
Impact Statement
Our work contributes a new method for modeling dynamical systems from unpaired observations with structural constraints. As such, our algorithm can contribute to biomedical domains, for example, by modeling drug perturbation responses in spatially resolved tumor tissues or by analyzing complex biological dynamics with unpaired datasets. These applications could positively impact biomedical research by enabling more data-efficient modeling of longitudinal biological processes and expediting hypothesis generation about treatment effects or underlying mechanisms. Future work should emphasize broad validation of our method using experimental or clinical ground truth, and the interpretation of results with domain experts to build appropriate trust in predictions.
@@ A.1 Flow Matching
Normalizing flows provide a parametric framework for characterizing transformations of a random variable into desired distributions (Papamakarios et al., 2021). These transformations can be realized through either finite (Rezende and Mohamed, 2015) or continuous compositions (Chen et al., 2018). The loss functions used in such formulations typically require computing Jacobians or integrating the flows at each forward pass, making them computationally expensive. Flow matching (FM) (Lipman et al., 2023; Albergo and Vanden-Eijnden, 2022; Liu et al., 2022) addresses this limitation by reducing the training of the velocity field to a regression problem, thereby making normalizing flows substantially more scalable. To ensure valid conditional paths at intermediate time points, samples are coupled either randomly or via optimal transport (Pooladian et al., 2023; Tong et al., 2024). Owing to this scalability, FM has been rapidly adopted across scientific domains, including biology and the life sciences (Li et al., 2025). In transcriptomics, for example, Klein et al. (2024) employed an FM backbone to approximate OT maps for drug response modeling and cross-modal translation tasks. Entropic OT formulations have also been applied to infer cellular trajectories (Tong et al., 2024; Rohbeck et al., 2025), generate imaging-based cell morphology changes (Zhang et al., 2025), and simulate spatial transcriptomics data from histology images (Huang et al., 2025). Despite these advances, existing work does not address how to meaningfully incorporate biological prior knowledge to constrain the velocity field, limiting the biological plausibility of inferred trajectories.
@@ A.2 Optimal Transport
Omics studies frequently generate uncoupled measurements across conditions, modalities, or time points, which must be integrated into a unified representation to provide a more comprehensive view of the underlying biology. Optimal transport (OT) has recently gained popularity for this task, as it provides a geometry-based approach to couple probability distributions (Bunne et al., 2024; Klein et al., 2025). In spatial transcriptomics (ST), several OT formulations have been introduced depending on context. For instance, Zeira et al. (2022) and Liu et al. (2023) proposed PASTE and PASTE2 to align ST data from adjacent tissue slices, while DeST-OT (Halmos et al., 2025) integrates spatio-temporal slices by modeling cell growth and differentiation. Rahimi et al. (2024) developed DOT, a multi-objective OT framework for mapping features across scRNA-seq and spatially resolved assays, and Ceccarelli et al. (2025) introduced TOAST, a spatially regularized OT framework for slice alignment and annotation transfer. While these methods are primarily designed to align biological data across space, time, or modality, they do not address the problem of trajectory inference toward biologically plausible solutions, leveraging biological priors to constrain or bias the transport plan.
Appendix B Additional Definitions
B.1 Kantarovich’s OT Formulation
B.2 Evaluation Metrics
2-Wasserstein. The 2-Wasserstein distance ($\mathcal{W}_{2}$ between empirical distributions $\mu,\nu$ is defined as:
where $\Pi(\mu,\nu)$ denotes the set of couplings between $\mu$ and $\nu$.
Weighted 2-Wasserstein. Implausible velocity fields can steer a cell’s transcriptional trajectory in unrealistic directions, potentially leading to entirely different terminal cell types. We thus employ the weighted 2-Wasserstein metric, which ensures the evaluation accounts for both transcriptional similarity and the distributional balance of cell types. We define the weighted 2-Wasserstein distance (Weighted $\mathcal{W}_{2}$) between true and predicted distributions as:
where $n_{i}^{\text{true}},n_{i}^{\text{pred}}$ are the number of true and predicted cells of type $i$, and $N$ is the total number of samples. To determine the cell type of generated trajectories, we employ a multi-class classifier $M_{\phi}$, implemented as an XGBoost model (Chen and Guestrin, 2016) trained for each dataset.
Energy Distance. Let $\mu$ and $\nu$ be probability distributions with samples $X=\{\bm{x}_{i}\}_{i=1}^{m}\sim\mu$ and $Y=\{\bm{y}_{j}\}_{j=1}^{n}\sim\nu$. The squared empirical energy distance (Energy) is defined as:
Maximum Mean Discrepancy. For the same samples, the unbiased empirical estimate of the squared maximum mean discrepancy (MMD) with kernel $\kappa$ is defined as:
Appendix C Proofs of Main Theoretical Results
Proposition C.1.
@@ Appendix C Proofs of Main Theoretical Results
$\Pi^{*}$ $={\arg\!\min}_{\Pi\geq 0}\;\sum_{k,l}\Pi_{kl}C_{kl}+\epsilon\sum_{k,l}\Pi_{kl}(\log\!\left(\Pi_{kl}\right)-1).$
@@
From Proposition 3 C.3, we have:
@@
$\Pi^{*}$ $={\arg\!\min}_{\Pi\geq 0}\;\sum_{k,l}\Pi_{kl}C_{kl}+\epsilon\sum_{k,l}\Pi_{kl}(\log\!\left(\Pi_{kl}\right)-1).$
@@
$\Pi^{*}$ $={\arg\!\min}_{\Pi\geq 0}\;\sum_{k,l}\Pi_{kl}C_{kl}+\epsilon\sum_{k,l}\Pi_{kl}(\log\!\left(\Pi_{kl}\right)-1).$
@@
$\epsilon\log\left(\frac{\Pi^{*}_{kl}}{M_{kl}}\right)=f_{k}-C_{kl}+g_{l}$
@@
From equations 16 15, 17 16, and 18 17 above, we get:
@@
Sinkhorn Algorithm uses these updates, iteratively, and these updates are shown to converge in (Franklin and Lorenz, 1989). Thus, Sinkhorn Algorithm can be used for the ContextFlow’s Prior Aware Entropy Regularized (PAER) (CTF-H) formulation.
From equation (9) 15, we get:
@@
Appendix D Effects of normalization on Prior aware cost matrix
From (Peyré et al., 2019), we know that optimal MOTFM coupling takes the form $\Pi^{*}_{\text{EOT}}=\text{diag}(u)\cdot K\cdot\text{diag}(v)$, where $K$ is the kernel matrix such that $K]_{ij}=\exp(\frac{-c_{ij}}{\epsilon})$, with $u,v$ satisfying marginalization constraints $u\odot Kv=a$ and $K^{T}u\odot v=b$. Sinkhorn updates are given by:
@@ Appendix D Effects of normalization on Prior aware cost matrix
$\Pi^{*}$ $={\arg\!\min}_{\mathbf{\Pi}\geq 0}\;\sum_{k,l}\Pi_{kl}C_{kl}+\epsilon\sum_{k,l}\Pi_{kl}(\log\!\left(\Pi_{kl}\right)-1).$
@@
$\mathbf{\Pi}^{*}$ $={\arg\!\min}_{\mathbf{\Pi}\geq 0}\;\sum_{k,l}\Pi_{kl}C_{kl}+\epsilon\sum_{k,l}\Pi_{kl}(\log\!\left(\Pi_{kl}\right)-1)$
@@
Corollary C.2 can also be interpreted as supporting the results of Proposition C.1 and our intuition that normalizing has the same effect on the kernel matrix as increasing $\epsilon$, leading to more diffused couplings or couplings with increased entropy.
Appendix E Detailed Analysis of ContextFlow
E.1 Algorithm Pseudocode
@@
Algorithm 1 ContextFlow: Flow Matching with Spatial-Context-Aware OT Couplings
@@
1:Input: gene data $\{\mathbf{X}_{t_{1}},\cdots,\mathbf{X}_{t_{m+1}}\}$, spatial data $\{\mathbf{S}_{t_{1}},\ldots,\mathbf{S}_{t_{m+1}}\}$, parameters $\lambda$, $\alpha$, $\epsilon$, $\sigma$, $r$, $\eta$
2:Data-Preprocessing: Compute local neighborhood means using Nearest Neighbor Algorithm and Ligand-Receptor features $f_{\text{LR}}$ using LIANA+ $\triangleright$ As defined in Equation 9 and 10
3:Output: neural velocity vector field $u_{\theta}$
4:Initialize $\theta$
5:while training do
6:  for $i=1,2,\ldots,m$ do
7:   Sample a batch $\mathcal{B}=\{(\bm{x}_{i},\bm{x}_{i+1}):(\bm{x}_{i},\bm{x}_{i+1})\sim(\mathbf{X}_{t_{i}},\mathbf{X}_{t_{i+1}})\}$
8:   Construct TPM: $\mathbf{M}_{i,i+1}(\mathcal{B})\triangleright$ $\mathbf{M}_{i,i+1}$ is defined in Equation 11
9:   if “prior-aware cost matrix” then
10:     $C_{kl}\leftarrow\alpha\cdot\|\bm{x}_{i}(k)-\bm{x}_{i+1}(l)\|_{2}^{2}+(1-\alpha)\cdot[\mathbf{M}_{i,i+1}]_{kl}$ for any pair $(k,l)$
11:     $\mathbf{K}\leftarrow\exp(-\mathbf{C}/\epsilon)$
12:   else if “prior-aware entropy regularization” then
13:     $C_{kl}\leftarrow\|\bm{x}_{i}(k)-\bm{x}_{i+1}(l)\|_{2}^{2}$ for any pair $(k,l)$
@@
14:     $\mathbf{K}\leftarrow\widehat{\mathbf{M}}_{i,i+1}\odot\exp(-\mathbf{C}/\epsilon)$ $\triangleright$ $\widehat{\mathbf{M}}_{i,i+1}$ is defined in Equation 13
15:   end if
16:   Initialize $\bm{a}\leftarrow\frac{1}{n_{i}}\mathbf{1}_{n_{i}}$, $\bm{b}\leftarrow\frac{1}{n_{i+1}}\mathbf{1}_{n_{i+1}}$, $\bm{u}\leftarrow\mathbf{1}_{n_{i}}$, $\bm{v}\leftarrow\mathbf{1}_{n_{i+1}}$
17:   while not converged do
18:     $\bm{u}\leftarrow\bm{a}\oslash(\mathbf{K}\bm{v}),\>\bm{v}\leftarrow\bm{b}\oslash(\mathbf{K}^{\top}\bm{u})$ $\triangleright$ Run Sinkhorn algorithm
19:   end while